OpenStudy (anonymous):

The x in the problem is actually theta...so... show that 1+cot^2x - cos^2x-cos^2xcot2x=1

OpenStudy (anonymous):

\[1+\cot^2x-\cos^2x-\cos^2x \cot(2x)=1?\]

OpenStudy (anonymous):

oh i meant 1+cot^2x−cos^2x−cos^2xcot^2x=1?

OpenStudy (anonymous):

\[\begin{align*}1+\cot^2x-\cos^2x-\cos^2x\cot^2x&=1\\ 1-\cos^2x+\cot^2x-\cos^2x\cot^2x&=\\ 1-\cos^2x+\cot^2x\left(1-\cos^2x\right)&=\\ \left(1+\cot^2x\right)\left(1-\cos^2x\right)&= \end{align*}\] Any identities you can think of?

OpenStudy (anonymous):

the last one switch it to csc^2x(sin^2x) ?

OpenStudy (raden):

maybe like this : let's do from left side : 1+cot^2x - cos^2x - cos^2xcot2x = 1+cot^2x - (cos^2 x)(1+cot^2x) = (1+cot^2x)(1 - cos^2 x) = (1+cot^2x)sin^2 x = csc^2 x sin^2 x = 1/sin^2 x * sin^2 x = 1

OpenStudy (anonymous):

@kungfupichu, yes, that's it. Now what can you rewrite that as?

OpenStudy (anonymous):

Actually, @RadEn already showed what you can do.